Consider an asset that costs $665,000 and is depreciated straight-line to zero over its 8-year tax life. The asset is to be used in a 5-year project; at the end of the project, the asset can be sold for $141,000. If the relevant tax rate is 21 percent, what is the aftertax cash flow from the sale of this asset? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(665,000/8)=$83125

Hence book value as on date of sale=Cost-Accumulated depreciation

=665,000-(83125*5)

=249375

Hence loss on sale=249375-141,000=$108375

Hence  aftertax cash flow=Sale proceeds+(Tax rate*loss on sale)

=141,000+(0.21*108375)

=$163759(Approx).